A uniform rectangular board has a length of 8.0 cm and a width of 6 cm. Its center of 

gravity will be at a point along its diagonal ___ from vertex V


1
Expert's answer
2021-01-28T19:38:47-0500

As we know, the center of gravity of the rectangle lies where two diagonals intersect each other. Therefore, the coordinates of the point of the center of gravity of the uniform rectangular board can be found as follows:


"G=(\dfrac{Length}{2}, \dfrac{Width}{2})=(\dfrac{8.0\ cm}{2}, \dfrac{6.0\ cm}{2})=(4.0\ cm, 3.0\ cm)."